# New Online Communities: Graph Deep Learning on Anonymous Voting Networks to Identify Sybils in Polycentric Governance

### 13 Dec 2023 v2

@@ 3.1   Quantitative and Empirical DAO Research
Two further studies on MakerDAO reveal its importance to the decentralized governance of digital assets. Sun, Stasinakis, and Sermpinis [46] conducted a detailed investigation into voter coalitions within MakerDAO, utilizing a k-means clustering algorithm on 809 governance polls, uncovering the impact of voter coalitions on protocol stability and growth. Zhao et al. [56] bring an Organizational Management perspective to task management in MakerDAO, with a binary variable analysis of strategic versus operational voting among 378 proposals. Echoing others, their findings present challenges to decentralization due to the dominance of powerful stakeholders, a concern echoed across DAO studies.
@@
In an effort to understand and categorize the diversity of forms DAOs take, Wright [53] proposed methodologies for measuring DAO autonomy by drawing lessons from other autonomous systems. Axelsen, Jensen, and Ross [3] examined organizational typologies in DAOs to determine when a DAO can be deemed decentralized. Ziegler and Welpe [58] developed a taxonomy for DAOs by exploring 35 DAOs and used agglomerative clustering to categorize them into five clusters. Finally, Braun, Haeusle, and Karpischek [8] attempt to model DAOs using a Transaction Cost Economics (TCE) perspective and consider the impact of revenue sharing, incentives for staking (investing), decentralization, and a variety of voting schemes. Like others who considered the role of decentralization, they model decentralization in terms of an Herfindahl-Hirschman Index (HHI), which is formally similar to GINI calculations [45][47] and top-player [30] and whale analyses [23].
Herfindahl-Hirschman: HHI =
$\sum_{i=1}^{N}s_{i}^{2}$
@@ 6.2.1 Defining Sybils and Deanonymizing Social Networks
Prior research, like Backstrom, Dwork, and Kleinberg [4] tackled social network deanonymization, but (like Sweeney) used a dataset that was anonymized after-the-fact. The advantage of these approaches is that the method is supervised and has a clear objective. Unfortunately, such methods are not practical against real datasets that achieve their anonymity by cryptographically deriving new addresses. Their method fits their goal: they propose two graph-based techniques for deanonymization, an active one that requires the prior construction of a subgraph of attacking nodes (prior collusion) to reveal information about neighbours (by walking the graph or making strategic cuts), and a passive method that relies on graph isomorphisms. Narayanan and Shmatikov’s more recent work on social network deanonymization [34] draws inspiration from Backstrom et al., but used graph isomorphisms with auxiliary graphs to link anonymous identities with public identities, a more information rich context and a method that requires a small number of ‘seed’ nodes present in both graphs.
@@
The most effective research on deanonymizing Bitcoin transactions uses real labelled datasets and intuitive machine learning algorithms (Akcora et al. [1] and Paquet-Clouston et al. [38]) or graph deep learning (Zhou et al. [57]), but these studies still have significant limitations. Paquet-Clouston et al. [38] attempts to trace ransomware payments by constructing graphs from 7222 labelled seeds originating from a convenience sample of known ransomware addresses. They construct an address graph and a cluster graph, the latter uses the logic of sybils to combine nodes that have shared inputs (multiple-input clustering heuristics). However, their analysis is purely descriptive and does not attempt to make predictions. Akcora et al. [1] also investigate ransomware, use convenience sample of three datasets (which includes the Paquet-Clouston et al. dataset), and construct a transaction graph and an address graph. They use multiple heuristics, develop graph features, use DBSCAN and k-means algorithms to cluster addresses and a naive similarity search. Finally, Zhou et al. [57] distinguish between address clustering and account identification, the former we call ‘sybil identification’ and the latter ‘deanonymization.’ They construct a trimmed graph of Ethereum transactions, develop node features, and train a graph neural network (GNN) to search subgraph node embeddings using a hierarchical (weighted) graph attention autoencoder. As is typical, they use several pooling layers and non-linear activation layers (LeakyReLu). They use cross-entropy loss for the classifier. They improve the performance of the default GNN by using contrastive subgraph learning. This self-supervised methodology, while in a different context, is similar to the approach detailed below.
@@ 6.3   Experimental Sybil Identification Methodology
Our dataset covers governance voting for thousands of DAOs between 2020-07-18 and 2023-02-27 (959 days, 2.6 years), which covers approximately 7.2 million voting transactions. However, naive, exact counts of DAOs and their governance proposals are ineffective due to noise. For instance, upon manual inspection the dataset reveals many registered DAOs that failed to launch or were clearly set up as tests, and many governance proposals have durations far too short to be practical (see Figures 3 and 4).
@@
Figure 4: Histogram of filtered governance proposal durations.
A larger and more recent dataset is available, however due to the computational requirements of deep learning and the method’s broad effectiveness, we rarely analyzed the entire dataset. Instead, we found that fairly short timeframes (about one year) provided reasonable results and were computationally efficient.
@@
Variants of this ‘follow the money’ approach are typical in the research, but as we discovered experimentally, without exceptionally large graphs the likelihood of finding sybil chains is very low (even in poor k-anonymous datasets, like Bitcoin, Ethereum, and our Snapshot). For testing purposes, we include a random fake sybil in the dataset and identify it.
@@ 6.3.2 Graph Convolutional Neural Network (GCNN)
With the voting graph, we engineer learnable features and train a Graph Convolutional Neural Network (GCNN) to create vector embeddings for each node. Importantly, this method is not limited to extensive wallet re-use (which we do see), rather, because the method generates embeddings for all nodes, including singletons, it is potentially effective against all graphs, including k-anonymous ones.
There are any number of possible custom Graph Convolutional Neural Networks (GCNNs) tailored for node classification tasks on graphs, with a multitude of interchangeable layers for various tasks. For our method, the first few steps are relatively straightforward. First, construct a voting graph where known (labelled) voters and re-used wallets are combined into single nodes with multiple edges to proposal nodes (we use a multi- or hyper-graph to accommodate multiple edges from each node). Second, we engineer features to learn higher dimensional vector embeddings on all nodes. We use a simple reconstruction loss (Mean Squared Error or MSE) for training the GCNN. Third, we use the node embeddings to conduct a k-means vector clustering on unknown nodes. We then propagate the most common label associated with each cluster to relabel the actual, lower dimensional nodes to form the sybil cluster. Finally, we take the relabelled similarity graph and combine nodes with their labels in each cluster to form a higher dimensional, reduced-size graph. The resulting graph is an inductive prediction of sybil nodes.
@@
We develop features and construct a typical train/test/validation split for the forward pass of our GCNN. The first layer of the forward pass is fully connected to capture node features and to incorporate voting edges and their ‘power’ attributes; the second layer incorporates the edge and node features with a sequential multi-layer perceptron (MLP) with ReLu activation that is then passed into a convolutional layer with mean feature aggregation; the third layer utilizes temporal sequences that have been engineered for a Long Short Term Memory (LSTM) layer; and finally, the fourth layer uses a multi-headed Graph Attention Network (GAT) to search and better learn edge connectivity. Each of these layers has tuneable hyperparameters that were set by conducting a grid search. After each forward pass and backpropagation to set training weights, training loss (MSE) was calculated, visualized, and assessed. The resulting tensors are normalized on zero and were checked for sufficient numeric variability.
@@ 6.3.3 Fast Vector Embeddings Search using FAISS
In formal notation:
@@ 7   Results
Tested against the snapshot.org dataset, our method typically reduces the original voting graph by approximately 2-5% (subject to hyperparameters). The method produces a voting graph with real IDs that can be inspected for further forensic analysis or visualized (see Figures 8 and 9). Moreover, we can produce a few key sociometric measurements to begin the process of constructing community baselines for further research (see Table 2).
@@ 8   Future Research
To improve and further validate our inductive method of sybil identification we need benchmarks from other online communities. Moreover, more sophisticated and further development of machine learning features would enhance the method. Finally, much more computationally expensive ‘following the money’ analyses that trace coins further to their origins would add a valuable graph signal for supervised learning.
In addition to refining current methods, it is worth exploring other deep learning techniques. To adapt the method for detecting voting spam, we could instead detect anomalies within the node embeddings. Thus, instead of identifying similar nodes, this approach searches for outliers that could indicate spam activity. In coin voting contexts, where the integrity of each vote is tied to a digital token, distinguishing between voting anomalies (spam) and similarities (sybils) is critical for safeguarding the democratic processes within DAOs.

### 27 Dec 2023 v3

@@ 3.1   Quantitative and Empirical DAO Research
Two further studies on MakerDAO reveal its importance to the decentralized governance of digital assets. Sun, Stasinakis, and Sermpinis [47] conducted a detailed investigation into voter coalitions within MakerDAO, utilizing a k-means clustering algorithm on 809 governance proposals, uncovering the impact of voter coalitions on protocol stability and growth. Zhao et al. [57] bring an Organizational Management perspective to task management in MakerDAO, with a binary variable analysis of strategic versus operational voting among 378 proposals. Echoing others, their findings present challenges to decentralization due to the dominance of powerful stakeholders, a concern echoed across DAO studies.
@@
In an effort to understand and categorize the diversity of forms DAOs take, Wright [54] proposed methodologies for measuring DAO autonomy by drawing lessons from other autonomous systems. Axelsen, Jensen, and Ross [3] examined organizational typologies in DAOs to determine when a DAO can be deemed decentralized. Ziegler and Welpe [59] developed a taxonomy for DAOs by exploring 35 DAOs and used agglomerative clustering to categorize them into five clusters. Finally, Braun, Haeusle, and Karpischek [9] attempt to model DAOs using a Transaction Cost Economics (TCE) perspective and consider the impact of revenue sharing, incentives for staking (investing), decentralization, and a variety of voting schemes. Like others who considered the role of decentralization, they model decentralization in terms of an Herfindahl-Hirschman Index (HHI), which is formally similar to Gini calculations [46][48] and top-player [31] and whale analyses [24]. These calculations metaphorically determine the relative ‘market’ position of DAOs and their ‘user’ ranks.
Formally, we define the following:
Herfindahl-Hirschman Index:
$HHI=\sum_{i=1}^{N}s_{i}^{2}$
@@
We performed HHI (see Figure 1), Gini (see Figure 2) and $n$-top players (see Figure 3) analyses on our dataset of 7.2m votes.
Figure 1: Herfindahl-Hirschman Index
Figure 2: Gini Coefficients
Figure 3: n-top Players Analysis
Finally, Bellavitis, Fisch, and Momtaz [5] conduct time series analyses on 2,300 DAOs using a “DAO Analyzer” dataset and studied voting from 1,196 DAOs using a Snapshot dataset. In the “DAO Analyzer” dataset they analyzed the financial returns of investments in DAO tokens, highlighting a key misunderstanding about DAOs and the impact of coin voting. While it is common practice to purchase DAO governance tokens with the expectation that they will yield financial returns like a conventional cryptocurrency, they also are used for governance voting. Coin voting creates a duality of motivations which muddy the conceptual waters, leading us to wonder, when do coins circulate like an economy, return value like an investment, or permit sustainable resource extraction like a commons? While the ‘coin’ in ‘coin voting’ functions like money, we argue below that DAOs are better understood as a commons resource management problem where coin voting determines institutional change through polycentric governance.
In terms of governance, in the Snapshot dataset Bellavitis, Fisch, and Momtaz found that 15% of DAOs have a token ownership threshold to make governance proposals, 160 votes on average per voter, and the proposal approval rate has hovered around 90%. However, these are extremely naive counts that do not discount the considerable noise in DAO governance nor the obvious role of sybils, so while the results offer a helpful baseline, we question the accuracy of the findings. For instance, using slightly less naive counting methods on a much larger Snapshot dataset of 59,939 proposals, we find approximately 100 average votes per voter.
@@ 6.2.1 Defining Sybils and Deanonymizing Social Networks
Prior research, like Backstrom, Dwork, and Kleinberg [4] tackled social network deanonymization, but (like Sweeney) used a dataset that was anonymized after-the-fact. The advantage of these approaches is that the method is fully supervised and has a clear objective. Unfortunately, such methods are not practical against real datasets that achieve their anonymity by cryptographically deriving new addresses. Their method fits their goal: they propose two graph-based techniques for deanonymization, an active one that requires the prior construction of a subgraph of attacking nodes (prior collusion) to reveal information about neighbours (by walking the graph or making strategic cuts), and a passive method that relies on graph isomorphisms. Narayanan and Shmatikov’s more recent work on social network deanonymization [35] draws inspiration from Backstrom et al., but used graph isomorphisms with auxiliary graphs to link anonymous identities with public identities, a more information rich context and a method that requires a small number of ‘seed’ nodes present in both graphs.
@@
The most effective research on deanonymizing Bitcoin transactions uses real labelled datasets and intuitive machine learning algorithms (Akcora et al. [1] and Paquet-Clouston et al. [39]) or graph deep learning (Zhou et al. [58]), but these studies still have significant limitations. Paquet-Clouston et al. [39] attempts to trace ransomware payments by constructing graphs from 7222 labelled seeds originating from a convenience sample of known ransomware addresses. They construct an address graph and a cluster graph, the latter uses the logic of sybils to combine nodes that have shared inputs (multiple-input clustering heuristics). However, their analysis is purely descriptive and does not attempt to make predictions. Akcora et al. [1] also investigate ransomware, use convenience sample of three datasets (which includes the Paquet-Clouston et al. dataset), and construct a transaction graph and an address graph. They use multiple heuristics, develop graph features, use DBSCAN and k-means algorithms to cluster addresses and a naive similarity search. Finally, Zhou et al. [58] distinguish between address clustering and account identification, the former we call ‘sybil identification’ and the latter ‘deanonymization.’ They construct a trimmed graph of Ethereum transactions, develop node features, and train a graph neural network (GNN) to search subgraph node embeddings using a hierarchical (weighted) graph attention autoencoder. As is typical, they use several pooling layers and non-linear activation layers (LeakyReLu). They use cross-entropy loss for the classifier. They improve the performance of the default GNN by using contrastive subgraph learning. This self-supervised methodology, while in a different and somewhat easier context, is similar to the approach detailed below. The biggest difference is that they classified nodes directly using the GNN, whereas we use develop the embeddings first and then classify using a embeddings search. The second step is necessary to detect sybils because sybils are by definition unknown and with limited connectivity to other voter nodes. Thus we do not have labels for sybils and we cannot a priori determine the number classes for self-supervised learning. A straightforward classification task will not be able to identify sybils.
@@ 6.3   Experimental Sybil Identification Methodology
The complete dataset covers governance voting for thousands of DAOs between 2020-07-18 and 2023-02-27 (959 days, 2.6 years), which covers approximately 7.2 million voting transactions. However, naive counts of DAOs, voters, and their governance proposals are ineffective due to noise. For instance, upon manual inspection the dataset reveals many registered DAOs that failed to launch or were clearly set up as tests, and many governance proposals have durations far too short to be practical (see Figure 6).
@@
6.3.1 Graph Signal Filter Analysis
For our dataset of 7.2m votes, using the PyGSP Python library, we constructed several wavelet graph signal filters to better understand the frequency components of voter time series.
Exponential Window (Expwin) Filter
The Expwin filter, characterized by an exponential decay, is a low-pass filter with a frequency response expressed as:
$g_{\text{exp}}(\lambda)=e^{-\tau\lambda}$
where $\lambda$ are the eigenvalues of $L$, and $\tau$ controls the decay rate. This filter emphasizes the low-frequency components of the signal, ideal for denoising and highlighting global trends across the graph (see Figure 7).
Figure 7: Spectral Graph Exponential Window Transform Histogram.
Spectral Graph Wavelet Transform (SGWT)
Spectral Graph Wavelet Transform (SGWT) is used to analyze graph signals by computing the Laplacian operator $L$, which encodes the connectivity structure of the graph, and its eigenvalues and eigenvectors, analogous to the classical Fourier transform. When a signal $x$ on a graph undergoes SGWT using a filter $g_{i}$, it is decomposed into components corresponding to different frequency bands captured by the wavelet filters (see Figure 8). Mathematically, the transform is given by:
$c_{i}=g_{i}(L)x=Ug_{i}(\Lambda)U^{T}x$
where $c_{i}$ are the coefficients after applying the $i$-th filter, $U$ is the matrix of eigenvectors of $L$, $\Lambda$ is the diagonal matrix of eigenvalues of $L$, and $g_{i}(\Lambda)$ applies the filter in the spectral domain.
Figure 8: Spectral Graph Wavelet Transform Histogram.
Meyer Wavelet Filter Bank
The Meyer Wavelet filter bank provides multi-resolution analysis and is a tight frame, suitable for decomposing signals into different scales. Its general form in the spectral domain is:
$g_{\text{Meyer}}(\lambda)=\nu\left(\frac{\lambda}{\lambda_{\text{max}}}\right)$
where $\nu(\cdot)$ transitions smoothly between 0 and 1, typically using sine and cosine functions. It captures both global and local features, valuable for feature extraction and signal compression (see Figure 9).
Figure 9: Spectral Graph Wavelet Transform Histogram.
The Expwin filter is best for capturing overall smoothness, while the Meyer wavelet strikes a balance between smoothness and localization. The SGWT with various filters can provide a comprehensive understanding of the signal’s structure on the graph. High coefficients in the SGWT indicate regions where the signal strongly resonates with the filter, pointing to significant features at different frequency bands.
@@ 6.3.1 Graph Signal Filter Analysis
Variants of this ‘follow the money’ approach are typical in the research, but as we discovered experimentally, without constructing exceptionally large graphs the likelihood of finding sybil chains is very low (even in poor k-anonymous datasets, like Bitcoin, Ethereum, and Snapshot), because isomorphisms between induced subgraphs are rare in anonymous settings. For testing purposes, we include a random fake sybil in the dataset and identify it.
@@ 6.3.3 Graph Convolutional Neural Network (GCNN)
With the voting graph, we engineer learnable features and train a Graph Convolutional Neural Network (GCNN) to create vector embeddings for each node. Since there are any number of possible custom GCNNs tailored for node classification tasks on graphs, with a multitude of interchangeable layers for various tasks, we designed the mechanism inductively with iterative testing against the dataset.
For our method, the first few steps are relatively straightforward. First, construct a voting graph where known (labelled) voters and re-used wallets are combined into single nodes with multiple edges to proposal nodes (we use a multi- or hyper-graph to accommodate multiple edges from each node). Second, we engineer features to learn higher dimensional vector embeddings on all nodes. We use a simple reconstruction loss (Mean Squared Error or MSE) to capture node features for training the GCNN (we cannot use a more conventional classification training loss like cross-entropy because we do not know the sybil classes a priori; instead, we use the GCN only to create the embeddings). Third, we use the node embeddings to conduct a k-means vector clustering on unknown nodes. We then propagate the most common label associated with each cluster to relabel the actual, lower dimensional nodes to form the sybil cluster. Finally, we take the relabelled similarity graph and combine nodes with their labels in each cluster to form a higher dimensional, reduced-size graph. The resulting graph is an inductive prediction of sybil nodes.
@@
We develop features and construct a typical train/test/validation split for the forward pass of our GCNN. The first layer of the forward pass is fully connected, capturing node features and incorporating voting edges with their ‘power’ attributes. The second layer integrates edge and node features with a sequential multi-layer perceptron (MLP) using ReLu activation, which is then fed into a convolutional layer with mean feature aggregation. The third layer processes temporal sequences engineered for a Long Short-Term Memory (LSTM) layer. Finally, the fourth layer uses a multi-headed Graph Attention Network (GAT) to better learn and search edge connectivity. Each of these layers has tuneable hyperparameters that were set by conducting a grid search. After each forward pass and backpropagation to set training weights, training loss (MSE) was calculated, visualized, and assessed. The resulting tensors are normalized on zero and were checked for sufficient numeric variability.
@@ 6.3.4 Fast Vector Embeddings Search using FAISS
In formal notation, the complete method:
@@ 7   Results
Tested against the snapshot.org dataset, our method typically reduces the original voting graph by approximately 2-5% (subject to hyperparameters). The method produces a voting graph with real IDs that can be inspected for further forensic analysis or visualized (see Figures 13 and 14). Moreover, we can produce a few key sociometric measurements to begin the process of constructing community baselines for further research (see Table 2).
@@ 8   Future Research
To improve and further validate our inductive method of sybil identification we need benchmarks from other online communities. Moreover, more sophisticated and further development of machine learning features would enhance the method. Finally, despite the computational demands required to ‘follow the money,’ using this direct method to identify sybil chains is the most reliable signal we have, and if conducted at scale the method would provide an invaluable graph signal for supervised learning.
In addition to refining current methods, it is worth exploring other deep learning techniques. To adapt the method for detecting voting spam, we could instead detect anomalies within the node embeddings. Thus, instead of identifying similar nodes, this approach would search for outliers to indicate spam activity. In coin voting contexts, where the integrity of each vote is tied to a digital token, distinguishing between voting anomalies (spam) and similarities (sybils) is critical for safeguarding the democratic processes within DAOs.
Finally, deep learning methods can be exploited by malicious users. Making deep learning methods resilient to attack is an active area of research and a practical requirement for DAOs, which must protect their digital commons from theft and inappropriate extraction. Examples of attacks against GCNNs include adversarial attacks that manipulate input data to deceive models, model inversion attacks that reveal sensitive information, and poisoning attacks that corrupt training data. Developing robust models that can detect and resist these attacks is crucial. Additionally, understanding and mitigating the ethical implications of AI and ML in DAO governance, such as bias in decision-making processes, must be a priority. These challenges open up diverse research avenues, ranging from developing new algorithms for anomaly detection to ethical frameworks for AI deployment in decentralized governance structures.